d/dØ∫1/(1-cosxcosØ)dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
I = \frac{d}{d\varnothing }\int \frac{1}{1-cosxcos\varnothing }dx
Using Leibniz’s rule, we have
I = \int \frac{\partial }{\partial \varnothing }(\frac{1}{1-cosxcos\varnothing })dx
I = \int (\frac{-(cosxsin\varnothing) }{(1-cosxcos\varnothing)^{2} })dx
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty